### v2

# On the cohomology of finite-dimensional nilpotent groups and Lie rings [^1] Keywords: nilpotent group, nilpotent Lie algebra, cohomology, finite-dimensional theory MSC 2020 : 17B56, 03C60

Abstract
We give some vanishing results for the first classical cohomology group in the case of finite-dimensional nilpotent groups and Lie rings. Based on the established cohomological results, we derive some structural consequences : we prove a form of Frattini’s argument for definable connected Cartan subrings and we give a definable version of Maschke’s theorem for the action of a definable connected $p$-divisible abelian group.
1 Introduction
Under the impetus of Hochschild and Serre (see [8] and [9]), techniques from homological algebra have been successfully applied to the study of modules in the category of groups and of Lie algebras. For (finite) groups, cohomological methods play a prominent role in the classification of group extensions and had striking applications in number theory through Galois cohomology (see [10]). In the case of Lie algebras, a number of structural results for finite-dimensional (in the usual linear sense) solvable Lie algebras were obtained by Barnes by using cohomological tools (see [2] and [3]). The approach of Barnes is particulary relevant in our case; more precisely, in [2] he proved that the cohomology groups associated to a nilpotent Lie algebra of finite linear dimension vanishe if the submodule of invariants is trivial. The same result is true when one replaces nilpotent Lie algebras by nilpotent finite groups or algebraic groups, nevertheless, a general and uniform treatement for this kind of vanishing results seems to be lacking.
Model-theory could provide this kind of general framework and we generalize Barnes’ methods to the context of groups and Lie rings (Lie algebras devoided of their vector space structure) equipped with some kind of dimension function on the definable subsets, which generalizes the properties of the linear dimension function and cardinality. From a model-theoretic point of view, we will consider groups and Lie rings that are definable in a finite-dimensional theory (see below for a precise definition). Finite groups, algebraic groups over an algebraically closed field, semi-algebraic groups over the reals are some striking examples of such "finite-dimensional" groups. On the other hand, finite-dimensional Lie algebras are the main instances of "finite-dimensional" Lie rings. Nevertheless, we decided to focus on the abstract setting of Lie rings since this kind of structures have recently captured the attention of the model-theoretic community (see [5] and [6] for example) and we hope what is presented here will contribute to this line of research. As an additional upshot, we revisite some standard algebraic results such as Frattini-like arguments and Maschke-like theorems.
Before stating our results, we say some words regarding the scope of our approach to cohomology. We will use two main tools : 1) the exactness of the inflation-restriction sequence for $n=1$, which can be proved by using elementary computations on the cochains 2) the existence of "long" exact sequences of cohomology groups induced by short exact sequences of $G$-modules. In particular, the Hoshchild-Serre spectral sequence will play no role in our approach.
Regarding cohomology, we prove the following :
Theorem 1.1.
  1. 1.
    In a finite-dimensional theory, let $G$ be a definable connected nilpotent group and $A$ be a definable connected $G$-module. Suppose that $A^{G}=0$. Then $H^{1}(G,A)=0$.
  2. 2.
    In a finite-dimensional theory, let $G$ be a definable connected nilpotent group and $A$ be a definable connected $G$-module. Suppose that $A^{G}=0$. Then for all definable connected sub-modules $V\leq U$ contained in $A$, one has $(U/V)^{G}=0$.
We formulate our results for groups but we could have opted for the point of view of Lie rings. On the basis of our cohomological results, we give the two following algebraic applications:
Theorem 1.2.
  1. 1.
    (definable Maschke’s theorem) In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module $p$-elementary as an abelian group. Suppose that $A^{T}=0$. Then $A=A_{1}\oplus\dots\oplus A_{r}$, where each $A_{i}$ is a definable connected $T$-irreducible sub-module of $A$.
  2. 2.
    (Frattini’s argument for Lie rings) In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring, and let $\mathfrak{c}$ be a definable connected Cartan subring of a definable connected ideal $\mathfrak{i}$.Then
    $\mathfrak{g}=\mathfrak{i}+N_{\mathfrak{g}}(\mathfrak{c}).$
In the second section, we recall some basic results about the cohomology of groups and after reviewing the case of finite nilpotent groups, we give an elementary proof of the vanishing of the first cohomology group when the submodule of invariants is trivial for a definable connected nilpotent group in a finite-dimensional theory. In the third section, we give our first algebraic application and we prove a definable version of Maschke’s theorem for the action of a definable connected $p$-divisible abelian group. In the fourth section, we pass to the case of Lie rings and in particular, we give a version of Frattini’s argument for definable connected Cartan subrings.
We thank Adrien Deloro for his many comments and suggestions. We are also grateful to Moreno Invitti for his key observation regarding long exact cohomological sequences.
2 Cohomological results
2.1 Preliminaries on the cohomology of groups
We recall some basic facts regarding the cohomology of groups. Although the results presented in this subsection are fairly standard, we decided to provide complete elementary algebraic proofs that could easily generalize to the case of abstract Lie rings for which general cohomological results may fail. For general references on this topic, we refer to [8] and to [10].
Definition 2.1.
A $G$-module is the data of ($G$,AA), where $G$ is a group and $A$ is an abelian group, equipped with a group action $\cdot"$ of $G$ on $A$. Equivalently, this amounts to giving a group homomorphism $\rho$ between $G$ and $\operatorname{Aut}(A)\subseteq\operatorname{End}(A)$.
Definition 2.2.
Let $(G,A)$ be a $G$-module. We define a cochain complex as follows :
  • $C^{0}(G,A)=A$.
  • For every integer $n\geq 1$, $C^{n}(G,A)$ denotes the abelian group of functions between $G^{n}$ and $A$.
The differential $d_{n}$ between $C^{n}(G,A)$ and $C^{n+1}(G,A)$ is defined, for every map $f$ in $C^{n}(G,A)$, by:
$d_{n}f(g_{1},\dots,g_{n+1})$
$=g_{1}\cdot f(g_{2},\dots,g_{n+1})+(-1)^{n+1}f(g_{1},\dots,g_{n})$
$+\sum_{1\leq s\leq n}(-1)f(g_{1},\dots,g_{s}g_{s+1},\dots,g_{n+1})$
We have $d_{n+1}\circ d_{n}=0$. We set $C(G,A)=(C^{n}(G,A))_{n\in\mathbb{N}}$.
We may now define the associated cohomology groups :
Definition 2.3.
Given the complex defined above, we introduce for every integer $n$ the $n$-th cohomology group, denoted by $H^{n}(G,A)$ :
$H^{n}(G,A)=\ker(d_{n})/\operatorname{im}(d_{n-1}).$
The group $G$ acts on the cochain groups of the form $C^{n}(H,A)$, for $H$ a normal subgroup, in a way that is compatible with the differentials :
let $f\in C^{n}(H,A)$ and $x\in G$, then $(x\cdot f)(h_{1},\dots,h_{n})=x\cdot f(h_{1}^{x},\dots,h_{n}^{x})$.
For $n=0$ and $n=1$, it is possible to give a convenient algebraic description of the cohomology groups.
Fact 2.4.
  1. 1.
    $H^{0}(G,A)=A^{G}=\{a\in A:g\cdot a=a,\text{for all $g\in G$}\}$.
  2. 2.
    $H^{1}(G,A)=\operatorname{Der}(G,A)/\operatorname{IDer}(G,A)$, where
    $\operatorname{Der}(G,A)=\{f\in C^{1}(G,A):f(xy)=x\cdot f(y)+f(x)\}$
    and
    $\operatorname{IDer}(G,A)=\{f\in C^{1}(G,A):f(x)=x\cdot a-a\penalty 10000\ \text{for some $a$ in $A$}\}.$
    In particular, if $f\in\operatorname{Der}(G,A)$, then $f(1)=0$.
Remark 2.5.
We observe that $G$ acts trivially on $H^{1}(G,A)$: let $f]\in H^{1}(G,A)$, where $f$ is a derivation, and let $x\in G$; then
$(x\cdot f)(y)-f(y)=x\cdot f(y^{x})-f(y)$
$=x\cdot(x^{-1}\cdot f(yx)+f(x^{-1}))-f(y)$
$=f(yx)+f(1)-f(x)-f(y)=y\cdot f(x)+f(y)-f(x)-f(y)$
$=y\cdot f(x)-f(x)$
for all $y\in G$; in other words, $(x\cdot f)-f$ is an inner derivation and thus $x\cdot f]=[f]$.
We now turn to the question of the relations between exact sequences of $G$-modules and long exact sequence of cohomology groups.
Fact 2.6.
Every short exact sequence of $G$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces the following short exact sequence :
$0\rightarrow C^{1}(G,A)\overset{i^{*}}{\rightarrow}C^{1}(G,B)\overset{p^{*}}{\rightarrow}C^{1}(G,C),$
where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$. Moreover, $i^{*}$ and $p^{*}$ commute with the differential.
Proof.
Since $i$ and $p$ commute with the action of $G$, the formula for the differential allows us to show that $i^{*},p^{*}$ commute with it.
Let us show that $i^{*}$ is injective. Indeed, for $\phi\in C^{1}(G,A)$, $i^{*}(\phi)=0$ iff $i(\phi(x))=0$ for all $x\in G$ iff $\phi(x)=0$ for all $x\in G$ by injectivity of $i$.
Now, we turn to the equality $\operatorname{im}(i^{*})=\ker(p^{*})$. Let $\phi\in\operatorname{im}(i^{*})\subseteq C^{1}(G,B)$; then there exists $\psi\in C^{1}(G,A)$, such that $\phi=i\circ\psi$ and hence $p^{*}(\phi)=p\circ i\circ\psi$. Consequently, for every $x\in G$, $p(i(\psi(x))=0$ since $\operatorname{im}(i)\subseteq\ker(p)$, and thus $\phi\in\ker(p^{*})$.
Conversely, let $\phi\in\ker(p^{*})$; then $p(\phi(x))=0$ for every $x\in G$. Thus $\phi(x)\in\operatorname{im}(i)$, since $\operatorname{im}(i)=\ker(p)$. Therefore for every $x$ there exists an unique $y_{x}\in A$, such that $i(y_{x})=\phi(x)$ by injectivity of $i$. Now, it suffices to consider the following element of $C^{1}(G,A)$ : $\psi:x\mapsto y_{x}$. ∎
Corollary 2.7.
Every short exact sequence of $G$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces the following short exact sequence :
$0\rightarrow Der(G,A)\overset{i^{*}}{\rightarrow}Der(G,B)\overset{p^{*}}{\rightarrow}Der(G,C),$
where $i^{*},p^{*}$ are defined by $i^{*}(f)=i\circ f$ and $p^{*}(f)=p\circ f$.
Proof.
Since $i$ and $p$ are $G$-morphisms, the sequence is well-defined. Moreover, let $\phi\in\ker(p^{*})$ and consider the application $\psi:G\rightarrow A,x\mapsto y_{x}$, where $i(y_{x})=\phi(x)$. Then, for all $x,y\in G$, we have $i(x\cdot\psi(y)+\psi(x))=x\cdot i(\psi(y))+i(\psi(x))=x\cdot\phi(y)+\phi(x)=\phi(xy)=i(\psi(xy))$, so, by injectivity of $i$, $\psi(xy)=x\cdot\psi(y)+\psi(x)$. ∎
One important feature of cohomology is the possibility of getting a "long" exact sequence of cohomology groups from a short exact sequence of $G$-modules.
Fact 2.8.
Let $0\rightarrow A_{0}\rightarrow A_{1}\rightarrow A_{3}\rightarrow 0$ be a short exact sequence of $G$-modules. Then there exists a "long" exact sequence of cohomology groups of the following form :
$A_{0}^{G}\rightarrow A_{1}^{G}\rightarrow A_{2}^{G}$
$\overset{\delta_{0}}{\rightarrow}H^{1}(G,A_{0})\rightarrow H^{1}(G,A_{1})\rightarrow\ H^{1}(G,A_{2})$
The morphism $(\delta_{0})$ is called a connecting morphism.
Proof.
It suffices to apply the Snake Lemma to the following commutative diagram, where $\phi_{i}:A_{i}\rightarrow Der(G,A_{i}),a_{i}\mapsto(g\mapsto g\cdot a_{i}-a_{i}),i=0,1,2$ :
[…]
We now introduce the Hochschild-Serre inflation-restriction sequence.
Definition 2.9.
Let $(G,A)$ be a $G$-module and let $H$ be a normal subgroup of $G$.
  • For every $n\in\mathbb{N}$, we define the restriction map $\operatorname{res}:C^{n}(G,A)\longrightarrow C^{n}(H,A),f\mapsto f|_{H}$.
  • We define the inflation map $\inf:C^{n}(G/H,A^{H})\longrightarrow C^{n}(G,A),f\mapsto f\circ\pi$, where $\pi$ denotes the canonical projection between $G$ and $G/H$.
Since the inflation and restriction morphisms commute with the differentials, we can pass to the cohomology groups and we get a sequence $H^{1}(G/H,A^{H})\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)$. In fact, this sequence is exact, and it can be proved by an elementary algebraic argument.
Fact 2.10.
[10, Proposition 4, §6, chap. VII])
The sequence $0\rightarrow H^{1}(G/H,A^{H})\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)$ is exact.
Proof.
  • $\inf$ is injective : let $f\in\operatorname{Der}(G,A)$, such that $f$ is constant on classes modulo $H$ and such that for all $g\in G$, $f(g)=g\cdot a-a$ for some $a\in A$. Then for all $h\in H$, $0=f(1)=f(h)=h\cdot a-a$, so $h\cdot a=a$.
  • $\operatorname{im}(\inf)\subseteq\ker(\operatorname{res})$ : let $f\in\operatorname{Der}(G,A)$ be constant on classes modulo $H$; we have $f(1)=0=f(h)$ for all $h\in H$.
  • $\ker(\operatorname{res})\subseteq\operatorname{im}(\inf)$. Let $f]\in H^{1}(G,A)$, such that $f|_{H}]=[0]$ in $H^{1}(H,A)$; thus there exists $a\in A$ such that $f(h)=h\cdot a-a$ for all $h\in H$. Up to considering the $f$-cohomologue $G$-derivation $\hat{f}$ defined by $\hat{f}(g)=f(g)-(g\cdot a-a)$, we may suppose that $f|_{H}=0$. Hence $f(gh)=f(g)+g\cdot f(h)=f(g)$ and $f$ is constant on classes modulo $H$; so $f(g)=f(gh)=f(hg)=f(h)+h\cdot f(g)=h\cdot f(g)$ since $H$ is a normal subgroup.
We can refine the image of the restriction morphism.
Fact 2.11.
$H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)^{G/H}$.
Proof.
Let $\phi]\in H^{1}(H,A)$ such that $\phi]=[\psi|_{H}]$ for $\psi\in\operatorname{Der}(G,A)$. As $\psi$ is a $G$-derivation, we are done by the same calculations as of Remark 2.5. ∎
The following Lemma will be very useful in the following.
Lemma 2.12.
Let $(G,A)$ be a $G$-module and let $H$ be a central subgroup of $G$, such that $H\leq C_{G}(A)$. Suppose that $A^{G}=0$. Then $H^{1}(H,A)^{G/H}=0$ and $H^{1}(G/H,A)\simeq H^{1}(G,A)$.
Proof.
Since $A^{H}=A$, by Fact 2.10 and Fact 2.11, the sequence $0\rightarrow H^{1}(G/H,A)\overset{\inf}{\rightarrow}H^{1}(G,A)\overset{\operatorname{res}}{\rightarrow}H^{1}(H,A)^{G/H}$ is exact. Therefore, it suffices to prove that
$H^{1}(H,A)^{G/H}=0.$
Let $f\in\operatorname{Der}(H,A)$; suppose that for all $x\in G$, $x\cdot f]=[f]$, i.e., for all $x\in G$ and for all $h\in H$, we have $x\cdot f(h^{x})-f(h)=x\cdot f(h)-f(h)=h\cdot a_{x}-a_{x}=0$ for some $a_{x}\in A$, since $H$ acts trivially on $A$ and $H\leq Z(G)$. In other words, for every $x\in G$, the equality $x\cdot f(h)=f(h)$ holds, and thus $f(h)\in A^{G}=0$; finally, we obtain that $f=0$. ∎
2.2 The case of finite groups
In this section, we give a quite elementary proof of the vanishing of the first cohomology group when $A^{G}=0$ for a finite nilpotent group $G$ and a finite $G$-module AA. This seems to be a quite classical result in the cohomology of finite groups, but our proof does not make use of Tate’s cohomology and it is more suitable to generalization to a model-theoretic framework. We draw inspiration of the proof strategy used by Barnes in [2] for finite-dimensional Lie algebras.
Let us recall a basic but fundamental lemma in representation theory : Schur’s Lemma.
Fact 2.13.
Let $A$ be a simple $R$-module, i.e $A$ has no non-trivial proper submodule. Then $C_{\operatorname{End}(A)}(R)$ is a skew-field.
We begin by the abelian case.
Lemma 2.14.
Let $G$ be a finite abelian group and $A$ be a finite $G$-module. Suppose that $A^{G}=0$. Then $H^{1}(G,A)=0$.
Proof.
We proceed by induction on the length, $\lg$, of $A$ in the category of finite $G$-modules. Since $A$ is finite, the length is also finite. If $A=0$, we are done.
  • $\lg(A)=1$, i.e., $A$ is irreducible. First, we way suppose that $G$ acts faithfully by Lemma 2.12 and so we have an embedding $\rho:G\hookrightarrow\operatorname{End}(A)$. Consider $R$ the subring of $\operatorname{End}(A)$ generated by $G$ and notice that $A$ is a simple $R$-module. By Schur’s Lemma, $C_{\operatorname{End}(A)}(R)$ is a skew field containing $\rho(G)$; this is even a commutative field by Wedderburn’s theorem.
    Now, let $f:G\rightarrow A$ be a non-trivial $G$-derivation. Let $y\in G-\{1\}$, such that $f(y)\neq 0$; then $(\rho(y)-Id)$ is invertible. Let $a=(\rho(y)-Id)^{-1}\cdot f(y)$. For $x\in G$, by commutativity of $G$ and the definition of a derivation, we have $(\rho(y)-Id)\cdot f(x)=\rho(x)\cdot f(y)-f(y)$, thus,
    $f(x)=(\rho(y)-Id)^{-1}\cdot\rho(x)\cdot f(y)-(\rho(y)-Id)^{-1}\cdot f(y)=\rho(x)\cdot a-a$
    by commutativity again. This shows that $H^{1}(G,A)=0$.
  • $\lg(A)>1$. Then, there exists a proper non-trivial $G$-module $B$, such that $\lg(B)<\lg(A)$ and $\lg(A/B)<\lg(A)$. Note that by hypothesis $B^{G}=0$. By induction on the length, we may therefore assume that $H^{1}(G,B)=0$. The long exact cohomology sequence will now come into play. Indeed, we have a short exact sequence of $G$-modules $0\rightarrow B\rightarrow A\rightarrow A/B\rightarrow 0$ which, by virtue of Fact 2.8, induces the following long exact cohomology sequence :
    $0\rightarrow B^{G}\rightarrow A^{G}\rightarrow(A/B)^{G}$
    $\overset{\delta_{0}}{\rightarrow}H^{1}(G,B)\rightarrow H^{1}(G,A)\rightarrow H^{1}(G,A/B)$
    By hypothesis, $A^{G}=B^{G}=0$ and by induction hypothesis, $H^{1}(G,B)=0$, so the exactness of the sequence then gives us that $(A/B)^{G}=0$; we may therefore apply once more the induction hypothesis and use the exactness of the long cohomology sequence to establish that $H^{1}(G,A)=0$.
Corollary 2.15.
Let $G$ be a finite nilpotent group and $A$ be a finite $G$-module. Suppose that $A^{G}=0$ and $A$ is $G$-irreducible. Then $H^{1}(G,A)=0$.
Proof.
We proceed by induction on the nilpotency class of $G$. We may suppose that $G$ is not abelian; since $G$ is nilpotent, we have $Z(G)\neq 1$. Notice that $A^{Z(G)}$ is a $G$-module. By irreducibility of $A$, we are reduced to two cases.
  • $A^{Z(G)}=A$. Then Lemma 2.12 gives $H^{1}(Z(G),A)^{G/Z(G)}=0$; so, the inflation-restriction exact sequence entails $H^{1}(G/Z(G),A)\simeq H^{1}(G,A)$ and we are done by induction.
  • $A^{Z(G)}=0$. Since $H^{1}(Z(G),A)=0$ (abelian case) and $H^{1}(G/Z(G),0)=0$, we have $H^{1}(G,A)=0$ by considering again the inflation-restriction exact sequence.
Corollary 2.16.
Let $G$ be a finite nilpotent group and $A$ be a finite $G$-module. Suppose that $A^{G}=0$. Then $H^{1}(G,A)=0$.
Proof.
We proceed exactly in the same way as in the proof of Lemma 2.14. ∎
To extend the proof to the category of definable $G$-modules in a first order theory, we need two main ingredients : 1) a notion of irreducible module that allows inductive reasoning on the lenght 2) an appropriate version of Schur’s Lemma regarding this irreducible category. In the next subsection, we will see that finite-dimensional theories satisfy these requirements.
2.3 Finite-dimensional nilpotent groups
We will study group cohomology inside the framework of finite-dimensional theories.
Definition 2.17.
[13, Definition 1.1] A theory $T$ is said to be finite-dimensional if there exists a dimension function, $\dim$, defined on the collection of all interpretable sets in models of $T$, with values in $\mathbb{N}$, such that for a formula $\phi(x,y)$ and interpretable sets $X,Y$, the following properties are satisfied :
  • (Invariance) If $a\equiv a^{\prime}$, then $\dim(\phi(x,a))=\dim(\phi(x,a^{\prime}))$.
  • (Finiteness) $\dim(X)=0$ iff $X$ is finite.
  • (Union) $\dim(X\cup Y)=\max\{\dim(X),\dim(Y)\}$.
  • (Fibration) If $f:X\longrightarrow Y$ is an interpretable map, such that $\dim(f^{-1}(y))\geq d$ for all $y\in Y$, then $\dim(X)\geq\dim(Y)+d$.
  • (Low fibration) If $f:X\longrightarrow Y$ is an interpretable map, such that $\dim(f^{-1}(y))\leq d$ for all $y\in Y$, then $\dim(X)\leq\dim(Y)+d$.
The dimension extends to type-definable sets, i.e., bounded intersections of definable sets, by taking the infimum of the dimensions of the definable sets which appear in the intersection.
The theory of a group of finite Morley rank and of an o-minimal structure are examples of finite-dimensional theories (for o-minimal structures, we must restrict to the collection of definable sets and interpretable group quotients). This model-theoretic framework is suitable for inductive reasoning when working with groups.
Fact 2.18.
[13, Proposition 2.3] Let $H\leq G$ be two type-definable groups in a finite-dimensional theory. If $H$ has infinite index in $G$, then $\dim(H)<\dim(G)$.
A definable group is said to be connected if it has no definable subgroup of finite index.
Another remarkable property of finite-dimensional theories concerns the possibility of obtaining a definable version of Schur’s Lemma.
Fact 2.19.
(adapted from [4]) Let $T$ be a finite-dimensional theory. Let $V$ be a definable connected abelian group and $S\leq\operatorname{End}_{def}(V)$ an invariant, i.e., a bounded union of type-definable sets, and unbounded subring of definable endomorphisms. Suppose that $V$ is an irreducible $S$-module, i.e., every proper definable connected $S$-invariant subgroup is trivial. Then $T=C_{End_{def}(V)}(S)$ is a definable skew-field.
In the following, unless explicitly stated otherwise, an irreducible module will be understood in the sense indicated above, i.e., in reference to the definable connected category.
Lemma 2.20.
In a finite-dimensional theory, let $(G,A)$ be a definable $G$-module. Suppose that $G$ is connected and abelian, that $A$ is connected and that $A^{G}=0$. Then $H^{1}(G,A)=0$.
Proof.
First, according to Corollary 2.12, up to considering $G/C_{G}(A)$, we may assume that $G$ acts faithfully. The result is trivial for $A=0$. We proceed by induction on the length, $\lg$, of $A$ as a $G$-module (in the definable connected category).
  • First, suppose $\lg(A)=1$, i.e., $A$ is irreducible. We denote by $\rho:G\rightarrow\operatorname{Aut}_{def}(A)\subseteq\operatorname{End}_{def}(A)$ the action. By Fact 2.19, we can linearize : there exists a definable skew-field $\mathbb{K}$ such that $\rho(G)\leq\mathbb{K}$. Up to considering $Z(\mathbb{K})$, which is also definable infinite, we may assume that $\mathbb{K}$ is commutative. Now, the proof goes as in the Lemma 2.14.
  • We may therefore assume $\lg(A)>1$; then there exists a non-trivial definable connected $G$-module $B$ strictly contained in $A$ such that $\lg(B)<\lg(A)$ and $\lg(A/B)<\lg(A)$. Now, the proof goes as in the Lemma 2.14.
At this point, we must consider a subclass of finite-dimensional theories.
Definition 2.21.
A finite-dimensional theory is said to be finite-dimensional if it satisfies the following property :
Any definable group admits a connected component, i.e., a smallest definable subgroup of finite index.
The theory of an o-minimal structure and the theory of a group of finite Morley rank belong to this subclass. In a finite-dimensional theory, a definable irreducible $G$-module has no proper infinite definable submodule. Moreover, an infinite definable nilpotent group has infinite center (see [1, Lemma 5.1, I]).
Theorem 2.22.
In a finite-dimensional theory, let $(G,A)$ be a definable $G$-module. Suppose that $G$ is connected and nilpotent, that $A$ is connected, and that $A^{G}=0$. Then $H^{1}(G,A)=0$.
Proof.
We proceed by induction on the nilpotency class of $G$. By Lemma 2.20, we may suppose that $G$ is not abelian. Notice that $G/Z(G)^{\circ}$ and $Z(G)^{\circ}$ are infinite definable connected nilpotent groups of nilpotency class strictly inferior.
First, suppose that $A$ is irreducible (in the definable connected category) and consider the definable submodule $A^{Z(G)^{\circ}}$. Then we have two cases :
  • $A^{Z(G)^{\circ}}=A$. By Lemma 2.12, $H^{1}(G,A)\simeq H^{1}(G/Z(G)^{\circ},A)$, and we are done by induction on the nilpotency class.
  • $A^{Z(G)^{\circ}}$ is finite. In fact, $A^{Z(G)^{\circ}}=0$ since $A^{G}=0$ and $G$ is connected. Notice that $H^{1}(G/Z(G)^{\circ},A^{Z(G)^{\circ}})=0=H^{1}(Z(G)^{\circ},A)$ by Lemma 2.20 and we are done by Fact 2.10.
Now, we proceed by induction on the length of $A$ as a $G$-module exactly in the same way as in the proof of Lemma 2.20. ∎
We conclude this section with a corollary on composition factors.
Corollary 2.23.
In a finite-dimensional theory, let $G$ be a definable connected nilpotent group and let $A$ be a definable connected $G$-module. Suppose that $A^{G}=0$. Then for all definable connected sub-modules $V\leq U$ contained in $A$, we have $(U/V)^{G}=0$.
Proof.
Indeed, we observe that $U^{G}=0=V^{G}$; then it suffices to consider the long exact cohomology sequence :
$0\rightarrow V^{G}\rightarrow U^{G}\rightarrow(U/V)^{G}$
$\overset{\delta_{0}}{\rightarrow}H^{1}(G,V)$
Theorem 2.22 allows us to conclude. ∎
3 Definable Maschke’s theorem
In this section, we provide an application of the cohomological study of finite-dimensional nilpotent groups by proving a definable version of Maschke’s theorem. We will closely follow the argument for the case of ranked groups given by Tindzhogo Ntsiri in [11]. Nevertheless, even for ranked groups, the result presented here is new since we can weaken the hypothesis "decent torus" and replace it by "definable connected abelian p-divisible". We will need the following definability lemma :
Fact 3.1.
([13, Proposition 3.6], extracted from [4]) In a finite-dimensional theory, let $\mathbb{K}$ be a skew-field such that :
  • there is an upper bound on dimensions of type-definable subsets of $\mathbb{K}$.
  • $\mathbb{K}$ contains an invariant unbounded subset.
Then $\mathbb{K}$ is definable.
Proposition 3.2.
In a finite-dimensional theory, let $V$ be a definable connected $p$-elementary abelian group and let $R$ be an unbounded invariant commutative ring of definable endomorphisms of $V$. Suppose that $R^{p}=R$. Let $W$ be a definable connected irreducible sub-$R$-module, such that $V/W$ is also $R$-irreducible. Then $W$ admits a definable $R$-invariant complement.
Proof.
By considering the action of $R$ on $W$ and $V/W$ and by applying Fact 2.19, the rings $C_{\operatorname{End}_{def}(W)}(R)=K_{W}$ and $C_{\operatorname{End}_{def}(V/W)}(R)=K_{V/W}$ are in fact definable skew-fields. Therefore $\operatorname{ann}_{R}(W)$ and $\operatorname{ann}_{R}(V/W)$ are prime ideals.
We may suppose that $\operatorname{ann}_{R}(W)=\operatorname{ann}_{R}(V/W)$. Indeed, let $0\neq r\in\operatorname{ann}_{R}(W)-\operatorname{ann}(V/W)$. The definable connected subgroups $\operatorname{im}(r)^{\circ}$ and $W\subseteq\ker(r)^{\circ}$ are $R$-invariant since $R$ is commutative. Then $\ker(r)^{\circ}/W$ and $(\operatorname{im}(r)^{\circ}+W)/W$ are $R$-invariant, and by irreducibility of $V/W$, we have $V=\operatorname{im}(r)^{\circ}+W$. Moreover, $\dim(V)=\dim(\operatorname{im}(r)^{\circ})+\dim(\ker(r)^{\circ})=\dim(\operatorname{im}(r)^{\circ})+\dim(\ker(r)^{\circ})-\dim(W\cap\operatorname{im}(r)^{\circ})$; thus $\dim(W\cap\operatorname{im}(r)^{\circ})=0$ and $\operatorname{im}(r)^{\circ}\cap W$ is finite. By linearizing again, $\operatorname{im}(r)^{\circ}\cap W$ is in fact trivial.
Conversely, let $0\neq r\in\operatorname{ann}_{R}(V/W)-\operatorname{ann}_{R}(W)$. Then $\operatorname{im}(r)^{\circ}=W$ by hypothesis and irreducibility of $W$. But $\ker(r)^{\circ}\cap W$ is finite by irreducibility of $W$ and it is in fact trivial by linearizing again. Therefore $V=W\oplus\ker(r)^{\circ}$.
Now, let $x\in\operatorname{ann}_{R}(W)=\operatorname{ann}_{R}(V/W)=I$; since $R=R^{p}$, there exists $y\in R$ such that $x=y^{p}$. In the domain $R/I$, we have $\overline{y}^{p}=\overline{x}=0$, so $y\in I$. Then $y\cdot V\subseteq W$ and $y^{2}\cdot V=\{0\}$. Finally, $x=0$ and $R$ embeds in $K_{W}$ and in $K_{V/W}$; by considering $Z(K_{W})$ (respectively, $Z(K_{V/W})$), we may suppose that $K_{W}$ (respectively, $K_{V/W}$) is commutative. Therefore $R$ acts by (scalar)-automorphisms on $W$ and $V/W$.
Now, we may suppose that $R$ acts by automorphisms on $V$ itself. Indeed, let $0\neq r\in R$ and suppose $\ker(r)\neq 0$. Then by irreducibility of $W$, $\ker(r)\cap W=0$; so by irreducibility of $V/W$, $V=W\oplus\ker(r)$.
In $\operatorname{End}_{def}(V)$, we consider the subring $K=R\cdot R^{-1}$; this is a commutative invariant and unbounded field. Let $Z\subseteq K$ be a type-definable subset. For $0\neq v\in V$, the evaluation map $K\rightarrow V,k\mapsto r\cdot v$ is injective since $K$ acts by automorphisms; thus $\dim(Z)=\dim(Z\cdot v)\leq\dim(V)$. By Lemma 3.1, $K$ is in fact a definable field. As a consequence, $W$ is a $K$-subspace of finite linear dimension and admits a definable $R$-invariant complement. ∎
Corollary 3.3.
(compare with [11, Proposition 2.13]) In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module that is $p$-elementary as an abelian group. Suppose that $A^{T}=0$. Then every definable connected irreducible sub-module of $A$ admits a definable $T$-invariant complement.
Proof.
By Corollary 2.23, the hypothesis $V^{T}=0$ is still true for definable connected $T$-invariant subgroups and sections of $V$. We proceed by induction on the length of $V$ as a definable connected $T$-module. Suppose $V\neq 0$; let $W$ be a proper definable connected irreducible sub-module. By induction, we may assume that $0\neq V/W$ is irreducible.
Let $R$ be the subring of $\operatorname{End}_{def}(V)$ generated by $T$. Since $T$ is abelian and $T^{p}=T$, we have $R^{p}=R$. Moreover, $W$ and $V/W$ are $R$-irreducible. Then, we are done by Proposition 3.2. ∎
Corollary 3.4.
In a finite-dimensional theory, let $T$ be a definable connected abelian group with trivial $p$-torsion and $A$ be a definable connected $T$-module that is $p$-elementary as an abelian group. Suppose that $A^{T}=0$. Then $V=V_{1}\oplus\dots\oplus V_{r}$, where each $V_{i}$ is a definable connected $T$-irreducible module.
Proof.
We may suppose that $V$ is not irreducible; let $V_{1}$ be a proper non-trivial definable connected sub-module. We proceed by induction on the length of $V_{1}$. If $\lg(V_{1})=1$, we are done by corollary 3.3. We may suppose that $\lg(V_{1})>1$; so, there exists a proper non-trivial definable connected submodule $V_{2}$ such that $\lg(V_{2})<\lg(V_{1})$ and $\lg(V_{2}/V_{1})<\lg(V_{1})$, and we can conclude by induction. ∎
4 The case of Lie rings
4.1 Cohomological results
In [9], Hochshild and Serre defines a cohomology theory for Lie algebras that parallels the theory for groups. The definition can easily be adapted to the more general context of Lie rings; nevertheless, general cohomological results for Lie algebras may fail when passing to Lie rings. Recall that a Lie ring $(\mathfrak{g},+,[,])$ is an abelian group equipped with a bi-additive and anti-symmetric function, $[,]$, that satisfies the Jacobi identity.
Definition 4.1.
A $\mathfrak{g}$-module is the data of ($\mathfrak{g}$,AA), where $\mathfrak{g}$ is a Lie ring and $A$ is an abelian group, equipped with a Lie ring action $\cdot"$ of $\mathfrak{g}$ on $A$. Equivalently, this amounts to giving a Lie ring homomorphism $\rho$ between $\mathfrak{g}$ and $\operatorname{End}(A)$.
Definition 4.2.
Let $(\mathfrak{g},A)$ be a $\mathfrak{g}$-module. We define a cochain complex as follows :
  • $C^{0}(\mathfrak{g},A)=A$
  • For every integer $n\geq 1$, $C^{n}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g}^{n},A):f\text{ is alternating}\}$, where $\operatorname{Hom}(\mathfrak{g}^{n},A)$ denotes the set of applications between $\mathfrak{g}^{n}$ and $A$, which are $\mathbb{Z}$-linear in each variable (we say also $\mathbb{Z}$-multilinear). Recall that a $\mathbb{Z}$-multilinear application $f$ is alternating if $f(x_{1},\dots,x_{n})=0$ whenever two arguments among the $x_{i}$ are equal.
The differential $d_{n}$ between $C^{n}(\mathfrak{g},A)$ and $C^{n+1}(\mathfrak{g},A)$ is defined, for every map $f$ in $C^{n}(\mathfrak{g},A)$, by :
$d_{n}f(g_{1},\dots,g_{n+1})$
$=\sum_{1\leq s<t\leq n+1}(-1)^{s+t-1}f([g_{s},g_{t}],g_{1},\dots,\hat{g_{s}},\hat{g_{t}},\dots,g_{n+1})$
$+\sum_{1\leq s\leq n+1}(-1)^{s}g_{s}\cdot f(g_{1},\dots,\hat{g_{s}},\dots,g_{n+1})$
The Lie ring $\mathfrak{g}$ acts on the cochain groups in a way that is compatible with the differentials :
Let $f\in C^{n}(\mathfrak{g},A)$ and $x\in\mathfrak{g}$, we define $x\cdot f$ by the formula
$(x\cdot f)(y_{1},\dots,y_{n})=x\cdot f(y_{1},\dots,y_{n})-\sum_{1\leq i\leq n}f(y_{1},\dots,y_{i-1},[x,y_{i}],y_{i+1},\dots,y_{n}),$
for all $y_{1},\dots,y_{n}\in\mathfrak{g}$. This action of $\mathfrak{g}$ is still valid if we replace $\mathfrak{g}$ by an ideal $\mathfrak{h}$ in the cochain groups.
For $n=0$ and $n=1$, it is also possible to give a convenient algebraic description of the cohomology groups.
Fact 4.3.
  1. 1.
    $H^{0}(\mathfrak{g},A)=A^{\mathfrak{g}}=\{a\in A:g\cdot a=0,\text{for all $g\in\mathfrak{g}$}\}$.
  2. 2.
    $H^{1}(\mathfrak{g},A)=\operatorname{Der}(\mathfrak{g},A)/\operatorname{IDer}(\mathfrak{g},A)$, where
    $\operatorname{Der}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g},A):f([x,y])=x.f(y)-y.f(x)\}$
    and
    $\operatorname{IDer}(\mathfrak{g},A)=\{f\in\operatorname{Hom}(\mathfrak{g},A):f(x)=x.a\penalty 10000\ \text{for some $a$ in $A$}\}.$
All the statements that we proved regarding the cohomology of groups are still true when passing to Lie rings, and the proofs require minor adaptations. In particular, we have the following :
Fact 4.4.
  • Every short exact sequence of $\mathfrak{g}$-modules $0\rightarrow A\overset{i}{\rightarrow}B\overset{p}{\rightarrow}C\rightarrow 0$ induces a "long" exact sequence of cohomology groups of the following form :
    $A^{\mathfrak{g}}\rightarrow B^{\mathfrak{g}}\rightarrow C^{\mathfrak{g}}$
    $\overset{\delta_{0}}{\rightarrow}H^{1}(\mathfrak{g},A)\rightarrow H^{1}(\mathfrak{g},B)\rightarrow\ H^{1}(\mathfrak{g},C)$
  • The sequence $0\rightarrow H^{1}(\mathfrak{g}/\mathfrak{h},A^{\mathfrak{h}})\overset{\inf}{\rightarrow}H^{1}(\mathfrak{g},A)\overset{\operatorname{res}}{\rightarrow}H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}$ is exact.
Lemma 4.5.
Let $(\mathfrak{g},A)$ be a $\mathfrak{g}$-module and let $\mathfrak{h}$ be a central ideal of $\mathfrak{g}$, such that $\mathfrak{h}\leq C_{\mathfrak{g}}(A)$. Suppose that $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g}/\mathfrak{h},A)\simeq H^{1}(\mathfrak{g},A)$.
Proof.
Since $A^{\mathfrak{h}}=A$, by Fact 4.4, the sequence $0\rightarrow H^{1}(\mathfrak{g}/\mathfrak{h},A)\overset{\inf}{\rightarrow}H^{1}(\mathfrak{g},A)\overset{\operatorname{res}}{\rightarrow}H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}$ is exact. Therefore, it suffices to prove that $H^{1}(\mathfrak{h},A)^{\mathfrak{g}/\mathfrak{h}}=0$.
Let $f\in\operatorname{Der}(\mathfrak{h},A)$; suppose that for all $\overline{x}\in\mathfrak{g}/\mathfrak{h}$, $(\overline{x}.f)]=[0]$, i.e., for all $y\in\mathfrak{h}$, we have $x.f(y)-f([x,y])=x.f(y)=y.a_{x}=0$ for some $a_{x}\in A$, since $\mathfrak{h}$ acts trivially and $\mathfrak{h}\leq Z(\mathfrak{g})$. In other words, for every $x\in\mathfrak{g}$, the equality $x\cdot f(y)=0$ holds, and thus $f(y)\in A^{\mathfrak{g}}=0$; finally, we obtain that $f=0$. ∎
There is also a minor variation concerning the abelian case for irreducible module.
Lemma 4.6.
In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A$ is $\mathfrak{g}$-irreducible and $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g},A)=0$.
Proof.
First, we way suppose that $\mathfrak{g}$ acts faithfully by Lemma 4.5 and so we have an embedding $\rho:\mathfrak{g}\hookrightarrow\operatorname{End}_{def}(A)$. Consider $R$ the subring of $\operatorname{End}_{def}(A)$ generated by $\mathfrak{g}$; by Fact 2.19 , $C_{\operatorname{End}_{def}(A)}(R)$ is a definable skew-field containing $R$; up-to considering its center, we may suppose that it is commutative.
Now, let $\delta:\mathfrak{g}\rightarrow A$ be a $\mathfrak{g}$-derivation. Let $y\in\mathfrak{g}-\{0\}$; then $\rho(y)$ is invertible. Let $a=\rho(y)^{-1}\cdot\delta(y)$. For $x\in\mathfrak{g}$, by commutativity and the definition of a $\mathfrak{g}$-derivation, we have :
$\rho(x)\cdot a=\rho(x)\cdot\rho(y)^{-1}\cdot\delta(y)=\rho(y)^{-1}\cdot\rho(x)\cdot\delta(y)=\rho(y)^{-1}\cdot\rho(y)\cdot\delta(x)=\delta(x).$
Thus $\delta(x)=\delta_{a}(x)$, where $\delta_{a}(x)=\rho(x)\cdot a$ is the $\mathfrak{g}$-inner derivation associated with $a$. This shows that $H^{1}(\mathfrak{g},A)=0$. ∎
Finally, we have the following :
Theorem 4.7.
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A^{\mathfrak{g}}=0$. Then $H^{1}(\mathfrak{g},A)=0$.
  • In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected nilpotent Lie ring and $A$ be a definable connected $\mathfrak{g}$-module. Suppose that $A^{\mathfrak{g}}=0$. Then for all definable connected sub-modules $V\leq U$ contained in $A$, one has $(U/V)^{\mathfrak{g}}=0$.
4.2 Definable Maschke’s theorem for restricted Lie algebras
Regarding Maschke’s theorem, we must pass to Lie algebras since the abstract Lie ring structure seems to be insufficient to capture a good notion of $p$-divisibility/$p$-unipotency. First, we clarify what semi-simplicity means for Lie algebras in positive characteristic. We start with the algebra of endomorphisms of a finite-dimensional vector space, for which linear algebra provides a natural notion of semi-simple element. For the following developments, we refer to Chapter II of [7].
Let $V$ be a finite-dimensional vector space over a field $\mathbb{K}$ of characteristic $p>0$; consider the associative $\mathbb{K}$-algebra $\mathfrak{g}=\operatorname{End}(V)$. We aim to characterize semi-simple endomorphisms purely in terms of algebraic operations in $\mathfrak{g}$ without reference to the minimal polynomial. To this end, raising to the $p$-th power will play a key role; we begin with the following definition :
Definition 4.8.
An element $x$ of $\mathfrak{g}$ is algebraically semi-simple if
$x\in\langle\bigcup_{i\geq 1}x^{p^{i}}\rangle_{vect}.$
The following fact shows that the two notions of semi-simplicity coincide :
Fact 4.9.
[7, Proposition 3.3, chap. 2] An element $x$ of $\mathfrak{g}$ is algebraically semi-simple if and only if it induces a semi-simple endomorphism.
For a Lie algebra over a field of characteristic $p>0$, one needs an analogue of the $p$-power map; this leads to the notion of restricted Lie algebra :
Definition 4.10.
A Lie algebra $\mathfrak{g}$ over a field $\mathbb{K}$ of characteristic $p>0$ is restricted if there exists a map $p]:\mathfrak{g}\to\mathfrak{g}$, $x\mapsto x^{[p]}$, satisfying the following conditions :
  1. 1.
    $x^{[p]},y]=ad_{x}^{p}(y)$ for all $y\in\mathfrak{g}$.
  2. 2.
    $(\lambda x)^{[p]}=\lambda^{p}x^{[p]}$ for all $\lambda\in K$.
  3. 3.
    $(x+y)^{[p]}=x^{[p]}+y^{[p]}+\sum_{1\leq i\leq p-1}s_{i}(x,y)$, where $ad_{x\otimes X+y\otimes 1}^{p-1}(x\otimes 1)=\sum_{1\leq i\leq p-1}is_{i}(x,y)\otimes X^{i-1}$ in $\mathfrak{g}\otimes_{K}K[X]$, for all $x,y\in\mathfrak{g}$.
Moreover, if $\mathfrak{g}$ is abelian, then $s_{i}(x,y)=0$ for all $i\in\{1,\dots,p-1\}$ and all $x,y\in\mathfrak{g}$.
Note that an associative $\mathbb{K}$-algebra with its natural Lie algebra structure is restricted when $[p]$ is taken as the $p$-th power. In particular, the algebra of endomorphisms of a vector space over a field of positive characteristic and the universal enveloping algebra of a Lie algebra over a field of positive characteristic have a natural structure of restricted Lie algebras.
Definition 4.11.
Let $(\mathfrak{g},[p])$ be a restricted Lie algebra over a field of characteristic $p>0$. An element $x$ is semi-simple if
$x\in\langle\bigcup_{i\geq 1}x^{[p]^{i}}\rangle_{vect}.$
In a finite-dimensional theory, a definable connected restricted Lie algebra $(\mathfrak{t},[p])$ is a torus if it is abelian and $ker([p])=0$.
Remark 4.12.
In the linear finite-dimensional case, an element of a restricted Lie algebra over a perfect field of characteristic $p>0$ is semi-simple if and only if it belongs to an abelian subalgebra stable under $[p]$, such that $ker([p])=0$ in this subalgebra (see [7, Proposition 3.3 and Theorem 3.7]).
Now, we may use Proposition 3.2 combining with Theorem 4.7 to prove the following :
Corollary 4.13.
In a finite-dimensional theory, let $\mathfrak{t}$ be a torus and $V$ a definable connected $\mathfrak{t}$-module. Assume that $V$ is abelian $p$-elementary and $\rho(t^{[p]})(x)=\rho(t)^{p}(x)$ for all $t\in\mathfrak{t}$ and $x\in V$, where $\rho$ is the morphism of representation induced by the action of $\mathfrak{t}$. Suppose that $V^{\mathfrak{t}}=0$. Then $V=V_{1}\oplus\dots\oplus V_{r}$, where each $V_{i}$ is a $\mathfrak{t}$-irreducible definable connected sub-module.
4.3 Frattini’s argument for Lie rings
In the theory of finite groups, $p$-Sylow subgroups play a fundamental role and often allow to obtain structural results about the ambient group. In particular, Frattini’s argument states that, for a finite group $G$, a normal subgroup $H$, and a $p$-Sylow subgroup $S$ of $H$, we have the following :
$G=HN_{G}(S).$
The argument is based on conjugacy results for $p$-Sylow and thus generalizes poorly to the setting of Lie rings : even when one has a suitable analogue of $p$-Sylow subgroups, the "inner" maps are not automorphisms of the Lie ring structure, and passing through the exponential map only partially remedies this defect. However, the conclusion of Frattini’s argument remains valid for finite-dimensional (in the usual linear sense) Lie algebras when $p$-Sylow subgroups are replaced by Cartan subalgebras. The proof relies on the cohomological study of nilpotent Lie algebras (see [3, Theorem 2.1]).
Taking into account our cohomological results and following [3, Theorem 2.1], we are now able to prove an analogous result in the context of finite-dimensional theories.
Definition 4.14.
In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring and let $\mathfrak{c}$ be a definable connected nilpotent subring such that
$N_{\mathfrak{g}}(\mathfrak{c})=\mathfrak{c}.$
We then say that $\mathfrak{c}$ is a definable connected Cartan subring.
Proposition 4.15.
In a finite-dimensional theory, let $\mathfrak{g}$ be a definable connected Lie ring, and let $\mathfrak{c}$ be a definable connected Cartan subring of a definable connected ideal $\mathfrak{i}$. Then
$\mathfrak{g}=\mathfrak{i}+N_{\mathfrak{g}}(\mathfrak{c}).$
Proof.
We set $\mathfrak{n}=N_{\mathfrak{g}}(\mathfrak{c})$ and we consider the definable connected $\mathfrak{c}$-module $A=\mathfrak{g}/\mathfrak{n}$.
Let us prove that $A^{\mathfrak{c}}=0$ for the adjoint action. Indeed, let $\overline{a}\in A$ be such that $c\cdot\overline{a}=\overline{0}$ for all $c\in\mathfrak{c}$; then $c,a]\in\mathfrak{i}\cap\mathfrak{n}$. But $\mathfrak{i}$ is an ideal and $N_{\mathfrak{i}}(\mathfrak{c})=\mathfrak{c}$, so $c,a]\in\mathfrak{c}$ for all $c$, hence $\overline{a}=\overline{0}$.
Now, consider the $\mathfrak{c}$-module
$(\mathfrak{g}/\mathfrak{n})/((\mathfrak{i}+\mathfrak{n})/\mathfrak{n})\simeq\mathfrak{g}/(\mathfrak{i}+\mathfrak{n}).$
If $(\mathfrak{g}/\mathfrak{n})/((\mathfrak{i}+\mathfrak{n})/\mathfrak{n})\neq 0$, by Theorem 4.7, $\mathfrak{c}$ acts non-trivially, a contradiction since $\mathfrak{c}\leq\mathfrak{i}$ and $\mathfrak{i}$ is an ideal.
Some applications of Frattini’s argument will be given in [12] (in collaboration with Tindzhogo Ntsiri). Notably, we prove that the Frattini subring is nilpotent in the context of ranked connected soluble Lie rings.
A generalization of some parts of this paper is currently under investigation by Moreno Invitti.